if sinA=cosB find the valueof  tan (A+B)
anser plzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzz

Grade:12th pass

2 Answers

Prakhar Kapoor
6 years ago
sin A = cos B
sin A = sin (90 – B )
 so,
A = 90 – B
i.e. A + B =90
therefore,
tan( A + B ) = tan ( 90 ) =infinity
sachin
4 years ago
In triagle 
Sum of all angle i.e A+B+C=180         _ __(1)
 A.T.Q .  SinA=CosB
               SinA=Sin(90-B)
               A=90-B.  ==>. A+B=90 .          ____(2)
   From equation 1 and 2 
  Angle C =90 
 CosC= Cos90 =0.